OpenStudy (anonymous):

Which pair are solutions to the equation? 4xy + 8 = 36 A. (1, 7) and (4, 9) B. (1, 7) and (7, 1) C. (7, 1) and (3, 2) D. (4, 9) and (3, 2) @iGreen :P

OpenStudy (anonymous):

@iGreen ?

OpenStudy (anonymous):

@ToMuchSwag2Handle

OpenStudy (igreen):

(1, 7) and (7, 1) are solutions to \( 4xy + 8 = 36 \)

OpenStudy (igreen):

So your answer is B

OpenStudy (anonymous):

@iGreen are you sure?

OpenStudy (igreen):

Yep, I graphed 'em. https://www.desmos.com/calculator

OpenStudy (anonymous):

..

OpenStudy (igreen):

Plus: 4xy + 8 = 36 4(1)(7) + 8 = 36 4(7) + 8 = 36 28 + 8 = 36 36 = 36 --------------------- 4xy + 8 = 36 4(7)(1) + 8 = 36 28(1) + 8 = 36 28 + 8 = 36 36 = 36

OpenStudy (igreen):

That's a graphing calculator @PLZHELPMEHhH

OpenStudy (anonymous):

okay @iGreen I have ONE LAST question.. ill post it in a sec :/

OpenStudy (anonymous):

@iGreen was right.

OpenStudy (anonymous):

@PLZHELPMEHhH are you doing FLVS?